# BraVE: Offline Reinforcement Learning for Discrete Combinatorial Action Spaces

### v2

@@ Abstract
Offline reinforcement learning in high-dimensional, discrete action spaces is challenging due to the exponential scaling of the joint action space with the number of sub-actions and the complexity of modeling sub-action dependencies. Existing methods either exhaustively evaluate the action space, making them computationally infeasible, or factorize Q-values, failing to represent joint sub-action effects. We propose Branch Value Estimation (BraVE), a value-based method that uses tree-structured action traversal to evaluate a linear number of joint actions while preserving dependency structure. BraVE outperforms prior offline RL methods by up to 20×20$\times$20 × in environments with over four million actions.  [^1] Code is available at https://github.com/matthewlanders/BraVE
@@ 1 Introduction
Our experiments show that BraVE consistently outperforms state-of-the-art baselines across a suite of challenging offline RL tasks with combinatorial action spaces containing up to 4 million discrete actions. In high-dimensional environments with strong sub-action dependencies, BraVE improves average return by up to 20×20$\times$20 × over state-of-the-art offline RL methods. While baseline performance degrades with increasing sub-action dependencies or action space size, BraVE maintains stable performance.
@@ 2 Preliminaries
In offline RL, the agent learns from a static dataset of transitions =$\mathcal{B}=\{(s_{t},a_{t},r_{t},s_{t+1})^{i}\}_{i=0}^{Z}$ generated by, possibly, a mixture of policies collectively referred to as the behavior policy $\pi_{\beta}$.
@@ 3.2 Tree Sparsification
Figure 2: BraVE traversal in a 3-D binary action space (full tree shown bottom-right). Starting from the root $[0,0,0]$, the agent selects $\hat{a}^{\prime}_{1}=1$ since its branch value ($11$) exceeds both those of alternative children ($4$, $-1$) and the root’s Q-value ($8$). Traversal proceeds until reaching $[1,1,0]$, where the Q-value ($16$) exceeds the child’s branch value ($1$); a terminal condition. Masked values (--) are ignored.
@@ 3.3 Q-Guided Tree Traversal
To compute the greedy action $\hat{\mathbf{a}}^{\prime}=\arg\max_{\mathbf{a}^{\prime}}Q(s^{\prime},\mathbf{a}^{\prime};\theta^{-})$ required by Equation (1), BraVE traverses the action space tree using a neural network $f(\cdot,\cdot;\theta^{-})$ to guide decisions at each node. At a node $\mathbf{a}_{\text{node}}$, The network receives as input the current state $s$ and the $N$-dimensional action vector corresponding to $\mathbf{a}_{\text{node}}$. It outputs a scalar node Q-value $q=Q(s,\mathbf{a}_{\text{node}};\theta^{-})$ and a vector of branch values $\mathbf{v}\in\mathbb{R}^{M_{\text{children}}^{\max}}$, where each $v_{j}$ estimates the maximum Q-value obtainable in the subtree rooted at the $j$-th child. While $q$ reflects the value of selecting the current node’s complete action, the branch values $\mathbf{v}$ estimate the best values available in each of the child subtrees.
@@
Traversal begins at the root node. At each step, the network evaluates $(q,\mathbf{v})$ at the current node. If $q\geq\max_{j}v_{j}$, traversal terminates and the corresponding action $\mathbf{a}_{\text{node}}$ is returned as $\hat{\mathbf{a}}^{\prime}$. Otherwise, the algorithm proceeds to the child indexed by $\arg\max_{j}v_{j}$. This greedy descent continues until either the termination condition is satisfied or a leaf node is reached. Notably, traversal may terminate at internal nodes rather than always proceeding to a fully specified leaf. This mechanism approximates the action with the highest estimated Q-value over the full combinatorial space, while requiring only a linear number of evaluations. A full example of this traversal process is shown in Figure 2.
@@ 3.4 Loss Computation
as the supervisory signal. This target is then propagated recursively up the tree. At each step, the target is used to update the parent node’s estimate for the corresponding branch value. The propagated target is updated at each level to reflect the maximum of $Y$ and the existing sibling branch values, ensuring consistency with the max-based traversal logic. This process continues up to the root, training each node to estimate the highest Q-value accessible through its subtree. The full $L_{\text{BraVE}}$ computation procedure is provided in Algorithm 1 and illustrated in Figure 3.
@@
2:$f(\theta)$: neural network with parameters $\theta$
3:$f(\theta^{-})$: target network with parameters $\theta^{-}$
4:$\{s,\mathbf{a},r,s^{\prime},\mathbf{a}^{\prime}\}$: transition from $\mathcal{B}$
5:$\hat{\mathbf{a}}^{\prime}$: action selected via tree traversal given $s$
6:$(q,\mathbf{v})\leftarrow f(s,\mathbf{a};\theta)$
7:$(q^{\prime},\mathbf{v}^{\prime})\leftarrow f\left(s^{\prime},\hat{\mathbf{a}}^{\prime};\theta^{-}\right)$
8:$Y\leftarrow\lambda\left(r+\gamma q^{\prime}\right)-\|\hat{\mathbf{a}}^{\prime}-\mathbf{a}^{\prime}\|$
@@
[…]
@@
[…]
@@
24:return total loss$/d$
@@ 3.5 Enhancements for Stability and Policy Quality
We mitigate action selection errors due to inaccurate branch value estimates through two complementary mechanisms: a depth penalty applied during training and beam search used during inference. During training, we apply a depth-based weighting factor $\delta$ in the BraVE loss computation. Because the branch value target is propagated upward from a sampled action node to the root, this weighting amplifies the influence of errors closer to the root, where mistakes propagate to more of the tree. At inference time, we use beam search to improve action selection robustness. Instead of committing to a single greedy path, the algorithm retains the top-$W$ actions (ranked by predicted values) at each tree level. The final action is selected from the union of all beams, allowing for broader exploration of high-value combinations and improving policy quality.
@@ 4 Experimental Evaluation
The agent starts from a fixed origin $s_{0}$ and must reach a designated goal $g$. At each timestep, it receives a negative reward $r=-\rho(s,g)$ proportional to its Euclidean distance from the goal. Episodes terminate upon reaching the goal (with reward $+10$) or falling into a pit, a failure state that incurs a penalty of $r=-10\cdot\rho(s_{0},g)$. This penalty structure ensures that failure is strictly worse than any successful trajectory, even those that are long or indirect.
@@ 4.1 Performance Across Reward Structures and Sub-Action Dependencies
We measure performance across 10 randomly generated environments in each of two settings: the first includes non-factorizable reward structures without sub-action dependencies, and the second includes both non-factorizable rewards and interdependent sub-actions. For each task, we vary dimensionality from 2-D ($|\mathcal{A}|=16$) to 11-D ($|\mathcal{A}|>4$ million), where each additional dimension introduces two new sub-actions. Each environment has size 5 in every dimension. The agent starts in the top-left corner and must reach a goal in the bottom-right. Results are averaged over five random seeds.
@@
We begin with environments that contain no pits. In these settings, the transition dynamics are factorizable across sub-actions. However, the reward function cannot be linearly decomposed. As shown in Table 1, FAS performs comparably to BraVE in low-dimensional tasks. As dimensionality increases, however, the modeling bias introduced by FAS’s linear decomposition causes the learned Q-values to diverge from the true values, degrading its performance. BraVE, by contrast, remains stable across all action space sizes, as it evaluates complete actions rather than marginal components. IQL performs significantly worse than both BraVE and FAS across all action space sizes, suggesting that methods not explicitly designed for combinatorial actions struggle even when transitions are factorizable. Complete learning curves are provided in Appendix B.1.
@@
We next evaluate performance under increasing sub-action dependency induced by hazardous transitions. In an 8-D environment ($|\mathcal{A}|=65,536$), we vary the number of pits from 5% to 100% of the 6,561 interior (non-boundary) states, ensuring that there is always a feasible path to the goal along the boundary of the state space. As shown in Table 2, BraVE maintains stable performance across all pit densities. FAS, by contrast, degrades sharply once pits occupy half of the interior, and IQL declines steadily as dependency strength increases. These results highlight BraVE’s ability to handle tightly coupled decision-making, even when adverse interactions between sub-actions are common. Complete learning curves are provided in Appendix B.2.
@@ 4.3 Ablations and Hyperparameters
We conduct ablation studies to assess the impact of important design choices in BraVE. We first examine the depth penalty parameter $\delta$, which scales the contribution of branch value errors based on a node’s depth, assigning greater weight to updates near the root where misestimates affect larger portions of the tree. Next, we vary the loss weight $\alpha$ to study the trade-off between the behavior-regularized TD loss and the BraVE loss. To isolate the effect of BraVE’s tree structure and Q-guided traversal, we compare against a DQN-based variant trained using BraVE’s loss (Equation 4) but without tree-based action selection or branch value propagation. Finally, we evaluate how the beam width used during inference affects both policy quality and computational cost. Full results are provided in Appendix D.
@@ 6 Discussion and Conclusion
While BraVE demonstrates strong performance, several limitations and opportunities for future work remain. First, the current implementation relies on a fixed, predefined ordering of sub-action dimensions for tree construction. This ordering can influence performance, particularly when sub-action dependencies do not align with the chosen sequence. Future work could explore methods for learning or adapting this ordering. Second, although BraVE scales with the number of sub-action dimensions ($N$), its efficiency may degrade when individual sub-action dimensions ($|\mathcal{A}_{d}|$) have extremely high cardinality, as this increases the branching factor and the size of the output layer for branch values. Third, computing $L_{BraVE}$ involves backpropagating through multiple parent node evaluations per sampled action. While necessary for consistent value estimation, this adds overhead compared to standard Q-learning updates, especially in deep trees. Inference, however, remains efficient due to linear-time traversal. Next, like all offline RL methods, BraVE’s performance depends on the quality and coverage of the offline dataset. Although tree sparsification leverages available data effectively, generalizing to unseen action combinations not represented in the sparsified tree remains a challenge. Finally, our empirical validation on the CoNE benchmark, while designed to test key capabilities, could be extended to a broader range of combinatorial tasks to further assess generalizability.

### v3

@@ Abstract
Offline reinforcement learning in high-dimensional, discrete action spaces is challenging due to the exponential scaling of the joint action space with the number of sub-actions and the complexity of modeling sub-action dependencies. Existing methods either exhaustively evaluate the action space, making them computationally infeasible, or factorize Q-values, failing to represent joint sub-action effects. We propose Branch Value Estimation (BraVE), a value-based method that uses tree-structured action traversal to evaluate a linear number of joint actions while preserving dependency structure. BraVE outperforms prior offline RL methods by up to $20\times$ in environments with over four million actions.  [^1] Code is available at https://github.com/matthewlanders/BraVE
@@ 1 Introduction
Our experiments show that BraVE consistently outperforms state-of-the-art baselines across a suite of challenging offline RL tasks with combinatorial action spaces containing up to 4 million discrete actions. In high-dimensional environments with strong sub-action dependencies, BraVE improves average return by up to $20\times$ over state-of-the-art offline RL methods. While baseline performance degrades with increasing sub-action dependencies or action space size, BraVE maintains stable performance.
@@ 2 Preliminaries
In offline RL, the agent learns from a static dataset of transitions $\mathcal{B}=\{(s_{t},a_{t},r_{t},s_{t+1})^{i}\}_{i=0}^{Z}$ generated by, possibly, a mixture of policies collectively referred to as the behavior policy $\pi_{\beta}$.
@@ 3.2 Tree Sparsification
Figure 2: BraVE traversal in a 3-D binary action space (full tree shown bottom-right). Starting from the root $[0,0,0]$, the agent selects $\hat{a}^{\prime}_{1}=1$ since its branch value (1111) exceeds both those of alternative children (44, 1-1) and the root’s Q-value (88). Traversal proceeds until reaching $[1,1,0]$, where the Q-value (1616) exceeds the child’s branch value (11); a terminal condition. Masked values (-) are ignored.
@@ 3.3 Q-Guided Tree Traversal
To compute the greedy action $\hat{\mathbf{a}}^{\prime}=\arg\max_{\mathbf{a}^{\prime}}Q(s^{\prime},\mathbf{a}^{\prime};\theta^{-})$ required by Equation (1), BraVE traverses the action space tree using a neural network $f(\cdot,\cdot;\theta^{-})$ to guide decisions at each node. At a node $\mathbf{a}_{\text{node}}$, the network receives as input the current state $s$, which may be discrete or continuous, and the $N$-dimensional action vector corresponding to $\mathbf{a}_{\text{node}}$. It outputs a scalar node Q-value $q=Q(s,\mathbf{a}_{\text{node}};\theta^{-})$ and a vector of branch values $\mathbf{v}\in\mathbb{R}^{M_{\text{children}}^{\max}}$. While $q$ reflects the value of selecting the current node’s complete action, the branch values $\mathbf{v}$ estimate the maximum Q-value obtainable in the subtree rooted at the $j$-th child. The network produces both $q$ and $\mathbf{v}$ in a single forward pass, with the branch values $\mathbf{v}$ representing a learned estimate of each subtree’s potential.
Importantly, the network $f$ is a standard MLP requiring no specialized architecture. The tree is not encoded into the network input; instead, it is an external structure that organizes the $\arg\max$ computation.
@@
Traversal begins at the root node. At each step, the network evaluates $(q,\mathbf{v})$ at the current node. If $q\geq\max_{j}v_{j}$, traversal terminates and the corresponding action $\mathbf{a}_{\text{node}}$ is returned as $\hat{\mathbf{a}}^{\prime}$. Otherwise, the algorithm proceeds to the child indexed by $\arg\max_{j}v_{j}$. This greedy descent continues until either the termination condition is satisfied or a leaf node is reached. Notably, traversal may terminate at internal nodes rather than always proceeding to a fully specified leaf. A full example of this process is shown in Figure 2.
This traversal mechanism enables BraVE’s computational efficiency: it evaluates a single node per sub-action dimension, resulting in $\mathcal{O}(N)$ complexity — a factor of $\prod_{d}|A_{d}|/N$ fewer evaluations than exhaustive scoring. The tree serves as the structural scaffold that makes this targeted, linear-time traversal possible.
@@ 3.4 Loss Computation
as the supervisory signal. This target is then propagated recursively up the tree. At each step, the target is used to update the parent node’s estimate for the corresponding branch value. The propagated target is updated at each level to reflect the maximum of $Y$ and the existing sibling branch values, ensuring consistency with the max-based traversal logic. This process continues up to the root, training the branch values $\mathbf{v}$ at each node through direct supervision from $L_{\text{BraVE}}$ to estimate the highest Q-value accessible through its subtree. The full $L_{\text{BraVE}}$ computation procedure is provided in Algorithm 1 and illustrated in Figure 3.
The recursive mechanism underlying $L_{\text{BraVE}}$ stabilizes learning by propagating training signals not only to the node corresponding to the sampled action but also to every ancestor on its path. Consequently, internal nodes — including those not directly sampled — receive gradient updates. Because all branches share a single global Q-network, these updates generalize across similar states and actions, enabling BraVE to form reliable estimates even in rarely-sampled regions of the tree.
@@
[…]
@@
13:  parent $\leftarrow\textsc{GetParent}(\text{node})$
@@
15:  children $\leftarrow\textsc{GetChildren}(\text{parent})$
@@
24:return total loss/d/d
@@ 3.5 Enhancements for Stability and Policy Quality
We mitigate the sensitivity of BraVE’s hierarchical structure to inaccurate branch value estimates using two complementary mechanisms: a depth penalty applied during training and beam search used during inference. During training, we apply a depth-based weighting factor $\delta$ in the BraVE loss computation. Because the branch value target is propagated upward from a sampled action node to the root, this weighting amplifies the influence of errors closer to the root, where mistakes propagate to more of the tree. At inference time, we use beam search to improve action selection robustness. Instead of committing to a single greedy path, the algorithm retains the top-$W$ actions (ranked by predicted values) at each tree level. The final action is selected from the union of all beams, allowing for broader exploration of high-value combinations and improving policy quality.
@@ 4 Experimental Evaluation
The agent starts from a fixed origin $s_{0}$ and must reach a designated goal $g$. At each timestep, it receives a negative reward $r=-\rho(s,g)$ proportional to its Euclidean distance from the goal. Episodes terminate upon reaching the goal (with reward +10+10) or falling into a pit, a failure state that incurs a penalty of $r=-10\cdot\rho(s_{0},g)$. This penalty structure ensures that failure is strictly worse than any successful trajectory, even those that are long or indirect.
@@ 4.1 Performance Across Reward Structures and Sub-Action Dependencies
We measure performance across 10 randomly generated environments in each of two settings: the first includes non-factorizable reward structures without sub-action dependencies, and the second includes both non-factorizable rewards and dependent sub-actions. For each task, we vary dimensionality from 2-D ($|\mathcal{A}|=16$) to 11-D ($|\mathcal{A}|>4$ million), where each additional dimension introduces two new sub-actions. Each environment has size 5 in every dimension. The agent starts in the top-left corner and must reach a goal in the bottom-right. Results are averaged over five random seeds.
@@
We begin with environments that contain no pits. In these settings, the transition dynamics are factorizable across sub-actions. However, the reward function cannot be linearly decomposed. As shown in Table 1, FAS performs comparably to BraVE in low-dimensional tasks but degrades as dimensionality increases. Because the reward depends on the state induced by the full action vector, the FAS approach of assigning rewards to individual sub-actions introduces a modeling bias that grows with action dimensionality, causing its learned Q-values to diverge from the true values. BraVE avoids this instability by evaluating complete joint actions directly, maintaining high performance across all action-space sizes. IQL performs significantly worse than BraVE across all action space sizes, suggesting that methods not explicitly designed for combinatorial actions struggle even when transitions are factorizable. Complete learning curves are provided in Appendix B.1.
@@
We next evaluate performance under increasing sub-action dependency induced by hazardous transitions. In an 8-D environment ($|\mathcal{A}|=65,536$), we vary the number of pits from 5% to 100% of the 6,561 interior (non-boundary) states, ensuring that there is always a feasible path to the goal along the boundary of the state space (where pits are never placed). As shown in Table 2, BraVE maintains stable performance across all pit densities. FAS, by contrast, degrades sharply once pits occupy half of the interior, and IQL declines steadily as dependency strength increases. These results highlight BraVE’s ability to handle tightly coupled decision-making, even when adverse interactions between sub-actions are common. These results are further contextualized by the normalized episode-length score reported in Appendix B.2, where complete learning curves are also provided.
@@
4.2 Robustness to Action Order
BraVE’s tree structure is defined by a fixed sequence of sub-action dimensions. To assess sensitivity to action ordering, we trained BraVE, FAS, and IQL using five random permutations of this sequence in the 8-D, 50% pit CoNE setting, which is empirically the most challenging 8-D variant (see Table 2). Results averaged over these permutations are reported in Table 4.
These findings are consistent with those in Table 2, showing that BraVE outperforms FAS by an order of magnitude and is roughly twice as effective as IQL. This demonstrates that BraVE’s performance is robust to the ordering of sub-action dimensions and does not rely on a favorable structural arrangement.
BraVE FAS IQL
[…]
Table 4: Performance on the 8-D, 50% pit CoNE task, averaged over five random permutations of the sub-action dimensions. BraVE’s effectiveness is not dependent on a specific tree construction.
@@ 4.4 Ablations and Hyperparameters
We conduct ablation and hyperparameter studies to quantify the contribution of each of BraVE’s design choices. First, we examine the depth penalty $\delta$, which scales the loss by a node’s position in the tree. A small penalty ($\delta=1$) consistently yields the best results, confirming the importance of prioritizing accuracy near the root. Second, we vary the loss weight $\alpha$ to analyze the trade-off between the behavior-regularized TD loss and the BraVE loss. Performance remains stable across a range of $\alpha$ values, but omitting the TD term entirely ($\alpha=0$) leads to significant degradation, particularly in higher-dimensional tasks. Third, we isolate the contribution of BraVE’s tree structure by training a standard DQN baseline, which omits the hierarchical traversal and branch value propagation. This variant fails to learn a viable policy, confirming that BraVE’s structured decomposition is critical to its success. Fourth, we evaluate the effect of beam width on performance and inference time. A width of 10 is sufficient for strong performance, with larger widths offering negligible gains at minimal runtime cost, reflecting the approach’s practical efficiency. Full results are presented in Appendix D.
@@ 6 Discussion and Conclusion
While BraVE demonstrates strong performance, several limitations and opportunities for future work remain. First, the tree structure is defined by a fixed ordering of the sub-action dimensions. While Section 4.2 demonstrates robustness to this choice in CoNE, future work could explore methods for learning or adaptively selecting the dimension sequence. Second, while BraVE’s inference complexity is linear in the number of dimensions ($N$), the size of the network’s output layer and the tree’s branching factor grow with the sub-action cardinality ($|\mathcal{A}_{d}|$), which may impact performance on tasks with extremely high-cardinality discrete actions. Third, the recursive update for $L_{\text{BraVE}}$ requires backpropagation through all parent nodes of a sampled action, incurring a greater computational cost during training than a standard Q-learning update. Fourth, our empirical validation is confined to the synthetic CoNE benchmark; future work should assess BraVE’s generalizability across a broader range of combinatorial tasks and real-world applications. Finally, adapting the BraVE framework to online learning is an important direction for future work, requiring the development of new mechanisms to manage the exploration-exploitation trade-off within the structured action space.